OpenStudy (anonymous):

Medal.. :) Select the equations that are parallel and perpendicular to y = one over fourx + 7 and that pass through the point (-8, 9). parallel: y = one over fourx - 11 perpendicular: y = -4x + 23 parallel: y = -4x + 11 perpendicular: y = one over fourx - 4 parallel: y = 4x - 7 perpendicular: y = negative one over fourx - 7 parallel: y = one over fourx + 11 perpendicular: y = -4x - 23

OpenStudy (anonymous):

Parallel line have same slope as of the given line. Perpendicular line have negative reciprocal slope as of the given line

OpenStudy (anonymous):

Umm OK so choice C??

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

Nope. Your slope = 1/4 parallel slope = 1/4 perpendicular = -4

OpenStudy (anonymous):

Choice B?!

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

I don't get this at all

OpenStudy (anonymous):

Lets do it this way. standard formula of line => y =mx+c We are given y =(1/4)x + 7 thus slope of the line (m) = 1/4 now we are given a point (-8,9) we have to find the equation of line that is parallel and perpendicular to y =(1/4)x+7 and passes through (-8,9) we will use point slope form (y-y1) = m(x-x1) (x1,y1) = (-8,9) m(parallel)=1/4 m(perpendicular)=-4 so what will be the equation of line?

OpenStudy (anonymous):

Um negative 1/4 -11

OpenStudy (anonymous):

parallel or perpendicular?

OpenStudy (anonymous):

point slope form (y-y1) = m(x-x1) (x1,y1) = (-8,9) m(parallel)=1/4 m(perpendicular)=-4 Parallel line y-9 = (1/4)(x-(-8)) y-9 =(1/4)x +2 y = 1/4x +11 Perpendicular y - 9= -4 (x-(-8)) y -9 = -4x -32 y =-4x -23

OpenStudy (anonymous):

My answer is D . thank you!

OpenStudy (anonymous):

Anytime:)